# Stealing Trust: Unraveling Blind Message Attacks in Web3 Authentication

### v1

@@ 7.2. Evaluation of Web3AuthGuard
To ensure a comprehensive evaluation, we set up a local malicious website capable of sending signature requests to MetaMask and conducted three rounds of testing: In the first round, the malicious website directly used a second set of five messages from each website to send signature requests, testing whether Web3AuthGuard could detect blind message attacks. In the second round, the malicious website randomly embedded each test message into a message from another website before sending signature requests. This was to assess whether Web3AuthGuard could identify attacks in the presence of verification flaws. In the third round, we directly visited 25 websites and performed Web3 authentication to evaluate whether Web3AuthGuard generated false positives.
@@
Our evaluation shows that Web3AuthGuard is robust without any false positives or false negatives. In addition, Web3AuthGuard effectively handles server-side verification flaws (V3), and its fuzzy matching can detect attacks even when malicious websites embed others’ messages into their own. However, Web3AuthGuard is not entirely free of false positives or negatives. In two alerts, Web3AuthGuard might produce false positives (blind message attack) in cases where messages from two different websites, lacking domain and name fields, have the same message structures. Also, Web3AuthGuard is ineffective against the vulnerability unchecked fields (V2); this vulnerability permits attackers to alter the message content (message body) at will, rendering any user-side detection ineffective. In our testing of 25 cases, five exhibited this vulnerability, as detailed in the body column of Table 5.3. Consequently, we have manually classified these five cases as failures. In conclusion, Web3AuthGuard was effective in 20 of the 25 tested websites, detecting blind message attacks. Another limitation is that Web3AuthGuard requires multiple messages from the same website to extract a valid template when the website’s message contains variable fields. A practical solution is to preload templates of popular websites into the wallet.
@@ 8. Discussion
  • Verifiable Source. The client can easily verify the source of the message. Meanwhile, the message should be unique for each website and cannot be impersonated.
@@
To implement the above security policy, the new protocol should mandate the presence of domain and nonce security fields in the message. At the same time, the protocol should specify the message format so that the client and server can automatically parse these fields. During the Web3 authentication process, the client (wallet) can automatically check these security fields in the message and prompt the user of potential security risks. During the verification process, except to verify the validity of the signature, the server also parses the fields in the message and verifies them one by one to ensure the integrity of the message. Although the end-to-end protocol can effectively solve the flaws in the current Web3 authentication protocol, it is difficult to implement in the short term because it requires upgrading both the server and the wallet.
2) Unique Identity. The private key can generate multiple public keys, allowing users to use a unique key for Web3 authentication on each website, effectively eliminating blind message attacks. This also enhances user anonymity, as attackers cannot link a user’s public key to all the websites they have visited. This solution can be implemented by upgrading the wallet: when a user uses Web3 authentication, the wallet checks for an existing public key for the website and generates a new one if none exists. However, this approach isolates the user’s digital assets across different websites. This issue can be addressed with account abstraction (EIP-4337) (Buterin et al., 2021), although it requires additional transaction fees.
@@
Our study conducted experiments in the wild, which may raise ethical concerns. We critically analyzed our work using the ethics framework (Kohno et al., 2023) to assess the ethical considerations and potential risks associated with our study. During the experiment, we used randomly generated accounts from a wallet to detect vulnerabilities in Web3 authentication; no normal users were affected. The APIs and functions we tested are all public functions of the websites, and we actively notified the websites with vulnerabilities and offered suggestions for improvements. To demonstrate the attacks, we built non-public websites and used randomly generated accounts representing both attackers and victims. We conducted the demonstration locally to ensure that no online malicious websites were created to mislead or deceive users. We adhered to Responsible Disclosure, and because some websites had not yet responded, we submitted our security report to the CVE database. We understand that once vulnerabilities in decentralized services are disclosed, attackers can exploit them. Therefore, we will only release the code of Web3AuthGuard and not disclose the test scripts of websites.
@@ 9. Related Work
To the best of our knowledge, we are the first to perform an in-depth study of security in the Web3 authentication. Our research relates to Web authentication and distributed digital identity.
@@
Web2 Authentication security focuses on verifying user identities to prevent unauthorized access, which includes Password-based Authentication (Munyendo et al., 2022; Pal et al., 2019; Ma et al., 2014), Multi-Factor Authentication (MFA) (Golla et al., 2021; Reynolds et al., 2018; Krombholz et al., 2019), Single Sign-On (SSO) (Jannett et al., 2022; Ghasemisharif et al., 2018; Calzavara et al., 2018), and Fast IDentity Online 2 (FIDO2)  (Lassak et al., 2021; Hanzlik et al., 2023; Lyastani et al., 2020). Despite its security criticisms, password-based authentication remains prevalent due to its simplicity. Multi-Factor Authentication, particularly two-step verification, increases password-based authentication’s security (Krombholz et al., 2019) by requiring supplemental factors like a one-time password via SMS or email. Single Sign-On streamlines authentication by enabling a single credential’s use across numerous platforms. Common SSO standards include OpenID Connect (OIDC) (Mainka et al., 2017), OAuth 2.0 (Philippaerts et al., 2022), and SAML (Somorovsky et al., 2012). FIDO2, which uses hardware tokens, offers an attractive alternative to password-based authentication. Lyastani et al. (Lyastani et al., 2020) overview the popular user authentication technologies, and they particularly focus on the potential of FIDO2 as a standard for passwordless authentication.
Web2 authentication is vulnerable to common security threats like password guessing (Pasquini et al., 2021b; Xu et al., 2021; Pasquini et al., 2021a; Yu and Martin, 2022), session hijacking (Drakonakis et al., 2020; Ghasemisharif et al., 2022; Sudhodanan and Paverd, 2022), phishing (Lain et al., 2022)(Alhogail and Alsabih, 2021), man-in-the-middle (MitM) attacks (Birge-Lee et al., 2021)(Karapanos and Capkun, 2014), and credential stuffing (Pal et al., 2022)(Wang and Reiter, 2021). Web3 authentication, which is based on cryptography, effectively reduces threats such as password guessing. However, since Web3 commonly uses traditional methods like cookies after authentication, it may still be vulnerable to session hijacking. Blind message attacks pose a unique challenge to Web3 authentication, as they can automatically select target websites based on address and blockchain data, unlike phishing, which can only mimic a specific website. Moreover, they do not require interception of communication (MitM), but rather bypass authentication by stealing user identities.
@@
Decentralized Identifiers (DIDs) are user-generated identifiers that circumvent the need for a centralized registry  (Liu et al., 2020)(Liao et al., 2022). DIDs play a crucial role in the Web3 ecosystem, where through public key authentication technology, Web3 applications can easily verify user identities. Other techniques such as zero-knowledge proofs (Feige et al., 1988)(Yang and Li, 2020) also hold relevance in this field. MPCAuth (Tan et al., 2023) is a multi-factor authentication system for distributed-trust applications that addresses ease-of-use and privacy challenges. Ansaroudi et al.(Ansaroudi et al., 2023) systematically analyze multiple digital identity wallets. Korir et al.(Korir et al., 2022) conducted a study on decentralized identity wallets, revealing user misconceptions about DIDs.
@@
Sign-In with Ethereum (SIWE, EIP-4361)(Chang et al., 2021) requires the inclusion of specific fields such as domain, nonce in the message, and crypto wallets supporting the SIWE protocol will verify these fields. Our research complements SIWE; of the 27 websites we examined that use the SIWE protocol (#10, #20), one (#20) had a replay attack risk. Some EIPs (Ethereum Improvement Proposals) provide signed data and identity standards. EIP-191 (Swende and Johnson, 2016) and EIP-712 (Bloemen et al., 2017) improve message readability and complex data signing, respectively, while EIP-725 (Vogelsteller and Yasaka, 2017) suggests a multi-key control scheme for proxy smart contracts. However, they have not gained widespread adoption due to issues like distributed server incompatibility (SIWE) and transaction fees (EIP-725).
@@ 10. Conclusion
This paper investigates the security risks associated with Web3 authentication and introduces the concept of Blind Message Attack. This attack enables attackers to gain unauthorized access to user accounts. Meanwhile, we propose advanced attacks, namely Replay Attack and Blind Multi-Message Attack. Our dynamic detection tool, Web3AuthChecker, successfully identifies 22 out of 29 real-world deployments of Web3 authentication that are at risk of blind message attacks. To better alert users, we have developed Web3AuthGuard, a client-side solution to detect Blind Message Attacks in crypto wallets, and showed that it can alert users immediately of potential attacks. Web3AuthChecker and Web3AuthGuard will be open-source to facilitate future research.
@@ Acknowledgements.
We thank the anonymous reviewers and our anonymous shepherd for their comments and suggestions. This research is supported in part by an Ethereum Foundation Academic Grant.

### v4

@@ 7.2. Evaluation of Web3AuthGuard
To ensure a comprehensive evaluation, we set up a local malicious website capable of sending signature requests to MetaMask and conducted three rounds of testing: In the first round, the malicious website directly used a second set of five messages from each website to send signature requests, testing whether Web3AuthGuard could detect blind message attacks. In the second round, we assessed whether Web3AuthGuard could identify attacks in the presence of verification flaws; the malicious website randomly embedded each test message into a message from another website before sending signature requests. In the third round, we directly visited 25 websites and performed Web3 authentication to evaluate whether Web3AuthGuard generated false positives.
@@
Our evaluation shows that Web3AuthGuard is robust without false positives or negatives. In addition, Web3AuthGuard effectively handles server-side verification flaws (V3), and its fuzzy matching can detect attacks even when malicious websites embed others’ messages into their own. However, Web3AuthGuard is not entirely free of false positives or negatives. In two situations, Web3AuthGuard might produce false positives (Blind Message Attack) in cases where messages from two different websites, lacking domain and name fields, have the same message structures. Also, Web3AuthGuard is ineffective against the vulnerability unchecked fields (V2); this vulnerability allows attackers to alter the message content (message body) at will, rendering any user-side detection ineffective. In our testing of 25 cases, five exhibited this vulnerability, as detailed in the body column of Table 5.3. Consequently, we have manually classified these five cases as failures. In conclusion, Web3AuthGuard was effective in detecting Blind Message Attacks on 20 of the 25 tested websites. Another limitation is that Web3AuthGuard requires multiple messages from the same website to extract a valid template when the website’s message contains variable fields. A practical solution is to preload templates of popular websites into crypto wallets.
@@ 8. Discussion
  • Verifiable Source. The client can easily verify the source of the message. Meanwhile, each website’s message should be unique and cannot be impersonated.
@@
To implement the above security policy, the new protocol should mandate the presence of domain and nonce security fields in the message. At the same time, the protocol should specify the message format so that the client and server can automatically parse these fields. During the Web3 authentication process, the client (wallet) can automatically check these security fields in the message and prompt the user of potential security risks. During the verification process, except to verify the validity of the signature, the server also parses the fields in the message. It verifies them one by one to ensure the integrity of the message. Although the end-to-end protocol can effectively solve the flaws in the current Web3 authentication protocol, it is difficult to implement in the short term because it requires upgrading both the server and the wallet.
2) Unique Identity. The private key can generate multiple public keys, allowing users to use a unique key for Web3 authentication on each website, effectively eliminating blind message attacks. This also enhances user anonymity, as attackers cannot link a user’s public key to all their visited websites. This solution can be implemented by upgrading the wallet. When a user uses Web3 authentication, the wallet checks for an existing public key for the website and generates a new one if none exists. However, this approach isolates the user’s digital assets across different websites. This issue can be addressed with account abstraction (EIP-4337) (Buterin et al., 2021), although it requires additional transaction fees.
@@
Our study conducted experiments in the wild, which may raise ethical concerns. We critically analyzed our work using the ethics framework (Kohno et al., 2023) to assess our study’s ethical considerations and potential risks. During the experiment, we used randomly generated accounts from a wallet to detect vulnerabilities in Web3 authentication; no normal users were affected. The APIs and functions we tested are all public functions of the websites, and we actively notified the websites with vulnerabilities and offered suggestions for improvements. To demonstrate the attacks, we built non-public websites and used randomly generated accounts representing both attackers and victims. We conducted the demonstration locally to ensure that no malicious websites were created online to mislead or deceive users. We adhered to Responsible Disclosure, and because some websites had not yet responded, we submitted our security report to the CVE database. We understand that once vulnerabilities in decentralized services are disclosed, attackers can exploit them. Therefore, we will only release the code of Web3AuthGuard and not disclose the test scripts of websites.
@@ 9. Related Work
To the best of our knowledge, we are the first to perform an in-depth study of security in Web3 authentication. Our research relates to Web authentication and distributed digital identity.
@@
Web2 Authentication security focuses on verifying user identities to prevent unauthorized access, which includes Password-based Authentication (Munyendo et al., 2022; Pal et al., 2019; Ma et al., 2014), Multi-Factor Authentication (MFA) (Golla et al., 2021; Reynolds et al., 2018; Krombholz et al., 2019), Single Sign-On (SSO) (Jannett et al., 2022; Ghasemisharif et al., 2018; Calzavara et al., 2018), and Fast IDentity Online 2 (FIDO2)  (Lassak et al., 2021; Hanzlik et al., 2023; Lyastani et al., 2020). Despite its security criticisms, password-based authentication remains prevalent due to its simplicity. Multi-Factor Authentication, particularly two-step verification, increases password-based authentication’s security (Krombholz et al., 2019) by requiring supplemental factors like a one-time password via SMS or email. Single Sign-On streamlines authentication by enabling a single credential’s use across numerous platforms. Common SSO standards include OpenID Connect (OIDC) (Mainka et al., 2017), OAuth 2.0 (Philippaerts et al., 2022), and SAML (Somorovsky et al., 2012). FIDO2, which uses hardware tokens, offers an attractive alternative to password-based authentication. Lyastani et al. (Lyastani et al., 2020) provide an overview of user authentication technologies, focusing on the potential of FIDO2 for passwordless authentication.
Web2 authentication is vulnerable to common security threats like password guessing (Pasquini et al., 2021b; Xu et al., 2021; Pasquini et al., 2021a; Yu and Martin, 2022), session hijacking (Drakonakis et al., 2020; Ghasemisharif et al., 2022; Sudhodanan and Paverd, 2022), phishing (Lain et al., 2022)(Alhogail and Alsabih, 2021), man-in-the-middle (MitM) attacks (Birge-Lee et al., 2021)(Karapanos and Capkun, 2014), and credential stuffing (Pal et al., 2022)(Wang and Reiter, 2021). Web3 authentication, which is based on cryptography and effectively reduces threats such as password guessing, may still be vulnerable to session hijacking since it commonly uses traditional session management. Blind message attacks pose a unique challenge to Web3, as they can automatically select target websites based on address and blockchain data, unlike phishing, which can only mimic a specific website. Moreover, they do not require interception of communication (MitM), but rather bypass authentication by stealing user identities.
@@
Decentralized Identifiers (DIDs) are user-generated identifiers that circumvent the need for a centralized registry  (Liu et al., 2020)(Liao et al., 2022). DIDs play a crucial role in the Web3 ecosystem, where through public key authentication technology, Web3 applications can easily verify user identities. Other techniques, such as zero-knowledge proofs (Feige et al., 1988)(Yang and Li, 2020) also hold relevance in this field. MPCAuth (Tan et al., 2023) is a multi-factor authentication system for distributed-trust applications that addresses ease-of-use and privacy challenges. Ansaroudi et al.(Ansaroudi et al., 2023) systematically analyze multiple digital identity wallets. Korir et al.(Korir et al., 2022) conducted a study on decentralized identity wallets, revealing user misconceptions about DIDs.
@@
Sign-In with Ethereum (SIWE, EIP-4361)(Chang et al., 2021) requires the inclusion of specific fields such as domain, nonce in the message, and crypto wallets supporting the SIWE protocol will verify these fields. Our research complements SIWE; of the 27 websites we examined that use the SIWE protocol (#10, #20), one (#20) had a replay attack risk. Some EIPs (Ethereum Improvement Proposals) provide signed data and identity standards. EIP-191 (Swende and Johnson, 2016) and EIP-712 (Bloemen et al., 2017) improve message readability and complex data signing, respectively, while EIP-725 (Vogelsteller and Yasaka, 2017) suggests a multi-key control scheme for proxy smart contracts. However, they have yet to gain widespread adoption due to issues like distributed server incompatibility (SIWE) and transaction fees (EIP-725).
@@ 10. Conclusion
This paper investigates the security risks associated with Web3 authentication and introduces the concept of Blind Message Attack. This attack enables attackers to gain unauthorized access to user accounts. Meanwhile, we propose advanced attacks, namely Replay Attack and Blind Multi-Message Attack. Our dynamic detection tool, Web3AuthChecker, successfully identifies 22 out of 29 real-world deployments of Web3 authentication that are at risk of Blind Message Attacks. To better alert users, we have developed Web3AuthGuard, a client-side solution to detect Blind Message Attacks in crypto wallets, and showed that it can alert users immediately of potential attacks. Web3AuthChecker and Web3AuthGuard are open source to facilitate future research.
@@ Acknowledgements.
We thank the anonymous reviewers and our anonymous shepherd for their comments and suggestions. The authors from Shandong University were supported in part by Taishan Young Scholar Program of Shandong Province, China (Grant No. tsqn202211001) and Xiaomi Young Talents Program. This research was also supported by an Ethereum Foundation Academic Grant.